Minute of KAGRA commissioning meeting
- Date: 2018/10/9 11AM~12PM
- Attendance: Miyakawa, Kokeyama, Eleonora, Enomoto, Miyoki
- Zoom: Ushiba, Izumi, Yamamoto, Kimura, Yokozawa, Kozakai, Oshino
X arm alignment
- Keeping Alignment day by day.
- PR2 and PR3 moved.
- If PR2 height is aligned, it causes misalignment.
- Opening top caver of chamber may cause misalignment by touching Periscope.
- PR2 will be tuned today again.
- Currently both beam are reaching to end.
- Final alignment will be checked 10/10 in the night shift.
- Buffle PD will be installed soon, then we need high power input beam.
- 5W will be out in the PSL this week.
- Calibration line injection.
- They need rebooting PC once today or tomorrow.
AC power at racks
- SC3 was almost full.
- New driver seemed to be not working. Miyo-kun tried it in other network, the it worked.
- Trouble might be network traffic in kAGRA control.
- EPICS gateway has been implemented. Network traffic should be much smaller now (~1/10).
- Still not perfect. Needs more tuning.
- Next meeting: 10/9.